OpenStudy (love.kat123):

In 1987 a recording artist sold 14,000 albums. In 1998 that same artist sold 4,683 albums. Find the rate of change in record sales from 1987 to 1998. A. 1/847 years per album B. –847 albums per year C. 847 albums per year D. –847 years per album @Adjax

OpenStudy (love.kat123):

@texaschic101 @Mokeira

OpenStudy (dangerousjesse):

\[\frac{Change~in~record~sales}{Change~in~time}\]

OpenStudy (texaschic101):

slope (rate of change) = (y2 - y1) / (x2 - x1) (1987,14000)...x1 = 1987 and y1 = 14000 (1998, 4683)...x2 = 1998 and y2 = 4683 now we sub slope(rate of change) = (4683 - 14000) / (1998 - 1987) slope = - 9317/11 reduces to - 847/1 or just - 847 you might better get a second opinion

OpenStudy (love.kat123):

thanks

OpenStudy (dangerousjesse):

\[\frac{14683-14000}{1998-1987} = \frac{-19317 ~albums}{11 ~ years} = -847\]

OpenStudy (dangerousjesse):

I meant 9317, but yeah.. Texas is right :P